# On Continuous Optimization for Constraint Satisfaction Problems

### v1

@@ Abstract
Constraint satisfaction problems (CSPs) are fundamental in mathematics, physics, and theoretical computer science. While conflict-driven clause learning Boolean Satisfiability (SAT) solvers have achieved remarkable success and become the mainstream approach for Boolean satisfiability, recent advances show that modern continuous local search (CLS) solvers can achieve highly competitive results on certain classes of SAT problems. Motivated by these advances, we extend the CLS framework from Boolean SAT to general CSP with finite-domain variables and expressive constraints. We present FourierCSP, a continuous optimization framework that generalizes the Walsh-Fourier transform to CSP, allowing for transforming versatile constraints to compact multilinear polynomials, thereby avoiding the need for auxiliary variables and memory-intensive encodings. Our approach leverages efficient evaluation and differentiation of the objective via circuit-output probability and employs a projected gradient optimization method with theoretical guarantees. Empirical results on benchmark suites demonstrate that FourierCSP is scalable and competitive, significantly broadening the class of problems that can be efficiently solved by CLS techniques.
Introduction
Constraint Satisfaction Problems (CSPs) are foundational in mathematics, physics, and theoretical computer science. Given a set of variables with domains taken from finite sets, and a set of constraints on the variables, i.e., a CSP instance, solving the instance is to find an assignment of variables that satisfies all the constraints. Applications of CSP solving are in artificial intelligence (Baluta et al. 2019), database query (Kolaitis and Vardi 1998), operational research (Brailsford, Potts, and Smith 1999), statistical physics (Nagy et al. 2024), information theory (Golia, Juba, and Meel 2022), quantum computing (Vardi and Zhang 2022), and so on.
Most CSP solvers can be categorized into two types: systematic search and local search. Systematic methods are often built upon the Branch-and-Bound (BnB) framework, which systematically explores the solution space while pruning infeasible branches. Notably, SAT solving, driven by Conflict-Driven Clause Learning (CDCL), i.e., BnB with clause learning (Davis and Putnam 1960; Davis, Logemann, and Loveland 1962; Marques-Silva and Sakallah 1999), has shown remarkable engineering success, routinely handling industrial-scale problems involving millions of variables (Vardi 2014). Nevertheless, SAT solvers primarily handle Conjunctive Normal Form (CNF) formulas, necessitating Booleanizations (Huang 2008) and subsequent CNF encodings (Prestwich 2021), potentially leading to substantial increases in formula size and inefficiency.
Integer Linear Programming (ILP) provides a more general formulation than CNF clauses, permitting more compact CSP representations through linearization, and utilizes linear programming relaxations to prune the search space efficiently. However, ILP solvers generally exhibit weaker Boolean reasoning capabilities than SAT solvers (Elffers and Nordström 2018). Constraint programming (CP) offers a native approach, directly accepting constraints expressed via a broad range of predicates over diverse variable types (Krupke et al. 2024). The research community of CP has developed multiple open-source solvers (Schulte, Lagerkvist, and Tack 2006; Prud’homme and Fages 2022; Hebrard 2008; Kuchcinski and Szymanek 2013). However, they typically do not scale as efficiently as ILP solvers and are less effective at constrained optimization problems. The OR-Tools CP-SAT solver, a portfolio solver with SAT, CP, has effectively addressed many limitations of traditional CP solvers and emerged as a robust alternative to ILP solvers, demonstrating competitive performance across a wide range of problems (Perron, Didier, and Gay 2023).
Systematic search solvers are complete; they can prove unsatisfiability or optimality if provided sufficient computation time. In contrast, local search methods aim to discover feasible solutions efficiently without proving unsatisfiability. Local search methods can be categorized into discrete local search (DLS) and continuous local search (CLS). DLS iteratively changes the value of variables to maximize the number of satisfied constraints (Mitchell, Selman, and Leveque 1992; Selman et al. 1993). CLS, is initiated by UniSAT (Gu 1994) and followed by FourierSAT (Kyrillidis et al. 2020), GradSAT (Kyrillidis, Vardi, and Zhang 2021), and FastFourierSAT (Cen, Zhang, and Fong 2025). CLS-based SAT solvers have been recently actively studied as a novel line of SAT frameworks. These solvers transform SAT into a polynomial optimization problem over the real domain and apply gradient-based optimizers. The latest work on this line, named FastFourierSAT, reveals that CLS solvers can be highly parallelizable and amenable to GPU accelerators. Recent theoretical and engineering efforts help CLS solvers achieve state-of-the-art (SOTA) performance on certain benchmarks in SAT solving, e.g., parity learning with error problems (Crawford, Kearns, and Schapire 1994).
Contributions
Despite promising advancements, current CLS methods are restricted to Boolean domains, limiting applicability to problems with inherently finite, multi-valued variables. In this work, we extend the FourierSAT framework to general CSPs, overcoming the significant challenge of directly mapping finite-domain variable domains to continuous optimization domains. We introduce an orthonormal product Fourier basis, allowing each constraint to be represented as a linear combination of these bases. The constraints are then reduced to a continuous polynomial with randomized rounding, where the input represents the probability of an assignment, and the output corresponds to the expectation of the constraint being satisfied. The polynomial representation of a constraint, i.e., the Walsh-Fourier expansion (WFE), can have an exponential number of terms, and a naive implementation is memory- and time-inefficient. In this work, we show that the WFE can be represented by the circuit-output probability (COP) of a decision diagram, which can be efficiently evaluated and differentiated.
We implemented our approach as FourierCSP with projected gradient ascent, The benchmark instances include both constraint satisfaction and optimization problems, including task scheduling and graph coloring with hashing query problems. Our approach is accelerated with GPU parallelization and compared with SOTA SAT-based, ILP-based, and CP-based complete solvers. Despite the incompleteness, our solver obtained encouraging results on CSP solving. To the best of our knowledge, this paper is the first algorithmic work to introduce continuous optimization to CSP. This opens a new research direction, enabling the integration of future continuous optimization techniques and advanced parallel hardware.
Preliminaries
@@ Introduction
Walsh-Fourier Transform is a method for transforming a function with a finite domain into a multilinear polynomial. Given the Fourier bases, any function defined on a finite domain has an equivalent Fourier representation.
Definition 1 (Fourier Bases).
@@
Theorem 1 (Walsh-Fourier Expansion (O’Donnell 2014)).
@@
Theoretical Framework
In this section, we first recap the previous Continuous Local Search (CLS) for Boolean SAT solving. We then demonstrate how our proposed solver, FourierCSP, can extend the CLS framework to general constraint satisfaction problems. Proofs are delayed to the supplemental material.
Continuous Local Search
CLS-based solvers define a continuous objective function, the optima of which encode the solutions to the original problem defined on discrete variables. The CLS framework is described in Algorithm 1
@@
Definition 2 (Objective).
Given a formula $F=\bigwedge_{c\in{}C}c$, every constraint $c\in{}C$ can be transformed to a function by Theorem 1. We denote the corresponding expansion as $f_{c}$ and the objective function w.r.t. $F$ by
$\mathcal{F}=\sum_{c\in{}C}f_{c}.$
FourierSAT, FastFourierSAT, and GradSAT are three gradient-based CLS variants. They transform Boolean constraints to continuous functions, called WFE via the Walsh-Fourier transform (O’Donnell 2014). However, the WFE of $n$ variables has at most $2^{n}$ terms. FourierSAT shows that, when the Boolean constraints are symmetric, e.g., XOR, cardinality, not-all-equal constraints, they can be concisely represented by $n$ elementary symmetric polynomials (Kyrillidis et al. 2020). FastFourierSAT demonstrates that elementary symmetric polynomials can be computed by convolutions, which correspond to multiplications in the frequency domain (Cen, Zhang, and Fong 2025). When the constraint is not symmetric, e.g., Pseudo-Boolean constraint, GradSAT can concisely represent the WFE by the COP of a binary decision diagram (BDD) (Kyrillidis, Vardi, and Zhang 2021).
Limitations of Previous CLS Solvers
The entire lineage of CLS solvers, from the 1990s approaches (Gu 1994) to modern variants like FourierSAT, FastFourierSAT, and GradSAT, is exclusively designed for Boolean constraints. While they perform well on certain problem classes, these solvers often fail to scale effectively to large or complex CSP instances. This limitation becomes particularly acute for finite-domain problems, which require conversion to potentially larger Boolean formulas. This limitation fundamentally restricts their practical application for real-world problems that naturally require finite-domain representations.
FourierCSP
To address the above limitation, we propose FourierCSP, a CLS framework that directly supports finite-domain CSPs without requiring Booleanization. We start by applying randomized rounding to the generalized form of WFE in Theorem 1. With the randomized rounding, we can reduce a constraint satisfaction problem to a continuous optimization problem. Subsequently, we reveal that the objective function is equivalent to the circuit-output probability of a decision diagram. Hence, the computational complexity of objective evaluation, i.e., one of the key bottlenecks for continuous optimizers, scales with the size of the decision diagram. Moreover, we derive practical convergence bounds that gradient-based methods can achieve within our framework.
@@
Theorem 1 illustrates how to expand discrete functions to multilinear polynomials through Walsh-Fourier expansion. The resulting Fourier coefficients $\hat{f}(\alpha)$ provide valuable insights into analyzing the discrete function properties. However, the expanded function $f$ remains discrete and non-differentiable given the Fourier Bases in Definition 1, limiting its utility for gradient-based optimization.
To enable continuous optimization, we introduce a randomized rounding $\mathcal{R}$ that relaxes discrete variable assignments into a continuous domain. Consider a set of finite domain variables $X=\{x_{1},...,x_{n}\}$. The vector probability space of $x_{i}$ is a simplex denoted by
@@
Definition 3 (Randomized Rounding).
@@
Definition 4 (Probability Space).
We define a probability space on a discrete matrix w.r.t. to real point $P\in\tilde{D}$, denoted by $\mathcal{S}:D\to\tilde{D}$, as:
@@
This probability space enables us to define a continuous extension of any discrete function by taking its expectation over a randomized assignment.
Theorem 2 (Expectation).
@@
$\mathop{\mathbb{E}}_{X\sim\mathcal{S}_{P}}[\mathcal{F}(X)]=\sum_{c\in{}C}\sum_{\alpha}\left(\hat{f}_{c}(\alpha)\cdot\prod_{i=1}^{n}p_{i,\alpha_{i}}\right),$ (1)
where the expectation is a multilinear polynomial.
Hence, the satisfiability problem of a CSP can be reduced to a continuous optimization problem by Theorem 3.
Theorem 3 (Reduction).
@@
$\max_{P\in\tilde{D}}\left(\mathop{\mathbb{E}}_{X\sim\mathcal{S}_{P}}[\mathcal{F}(X)]\right)=|C|.$ (2)
Circuit-Output Probability
Theorem 3 reveals that CSP solving can be reduced to a continuous optimization problem. The performance of gradient-based optimization crucially depends on the efficiency of evaluating and differentiating the objective function. However, the complexity of evaluation and differentiation is in #P-hard, i.e., Eq. (1) has at most exponentially many terms, in general. Motivated by (Thornton and Nair 1994), we reformulate Eq. (1) to circuit-output probability (COP), which admits more tractable computation when a structured representation can encode the underlying constraint, i.e., decision diagrams.
Definition 5 (Circuit-Output Probability).
Let $f_{c}$ be a discrete function over a set $X$ of finite domain variables. Let $P$ be the variable input probabilities. The circuit-output probability problem of $f_{c}$ w.r.t $P$, denoted by $\texttt{COP}_{c}(P)$, is the probability of $f_{c}$ outputs $1$ given the value of each variable independently sampled from a multimodal distribution, i.e.,
$\texttt{COP}_{c}(P)=\sum_{X\in D}f_{c}(X)\prod_{i=1}^{n}p_{i,x_{i}}.$
Corollary 1.
$\mathop{\mathbb{E}}_{X\sim\mathcal{S}_{P}}[f_{c}(X)]=\texttt{COP}_{c}(P).$
Let $\mathcal{C}(P)=\sum_{c\in{}C}\texttt{COP}_{c}(P)$, we have
$\mathcal{C}(P)=\mathop{\mathbb{E}}_{X\sim\mathcal{S}_{P}}[\mathcal{F}(X)]$
Corollary 1 implies that it is sufficient to evaluate Eq. (1) by computing the COP of each constraint. However, the complexity of calculating the COP is still in #P-hard for general representations. Nevertheless, if a constraint is represented by a decision diagram $G(V,E)$, then the COP on it can be obtained in $\mathcal{O}(|E|)$, where $|E|$ is the number of edges in $G$. Prior work demonstrates that BDD can encode many Boolean constraints in compact size (Kyrillidis, Vardi, and Zhang 2021; Sasao, Fujita et al. 1996; Aavani 2011).
The multiple-valued extension of BDD is called multi-valued decision diagram (MDD) (Bryant 1995). The MDD can be versatile in representing many constraints through the translation from table constraints (Cheng and Yap 2010). Similar to BDD, two MDDs can be combined through an APPLY operation (Bryant 1986; Perez and Régin 2015), offering flexibility in encoding versatile constraints. The MDDs can be reduced by merging the nodes having the same set of outgoing neighbors associated with the same label (Cheng and Yap 2010; Perez and Régin 2015). Moreover, the size can be further reduced by sifting (Miller and Drechsler 2003). A compact-size decision diagram means that the COP on it can be evaluated and differentiated efficiently.
Lemma 1 (Top-Down Evaluation).
$m_{td}[v]=\mathbb{P}_{M,a}[v],$
$m_{td}[\texttt{true}]=\texttt{COP}_{c}(P_{a}),$
i.e., Algorithm 2 returns the COP of $G$ in time $O(|E|)$.
Algorithm 2 Top-Down Traversal of MDD
Output Circuit-output probability $\texttt{COP}_{c}(P)$.
1: Let $m_{td}[v]$ be the top-down message of each node $v\in{}V$, where the root node is initialized as 1, otherwise 0.
2:for $v\in{}\texttt{sorted}(V)$ do
3:  for $u=\mathrm{child}(v)$ do
4:   $m_{td}[u]+=p_{\mathrm{id}(v),\mathrm{label}(v,u)}\cdot m_{td}[v]$
5:  end for
6:end for
7:return $m_{td}[true]$
Lemma 2 (Bottom-Up Evaluation).
$m_{bu}[v]=\texttt{COP}_{c,v}(P),$
i.e., Algorithm 3 returns the COP of $G$ and in time $O(|E|)$.
Algorithm 3 Bottom-Up Traversal of MDD
Output Circuit-output probability $\texttt{COP}_{c}(P)$.
2:for $v\in{}\texttt{sorted}(V,\texttt{reverse=True})$ do
3:  for $u\in\mathrm{child}(v)$ do
4:   $i=\mathrm{id}(v)$
5:   $m_{bu}[v]+=p_{\mathrm{id}(v),\mathrm{label}(v,u)}\cdot m_{bu}[u]$
6:  end for
7:end for
8:return $m_{bu}[root]$
The main idea of differentiation is from probabilistic inference (Pearl 1988; Shafer and Shenoy 1990), which traverses the graph model top-down and bottom-up. The following theorem shows that the partial derivative of COP can be obtained by top-down and bottom-up messages from Algorithm 2 and 3.
Theorem 4 (Derivatives).
$\frac{\partial\texttt{COP}_{c}(P)}{\partial p_{i,j}}=\sum_{v\in\{u:id(u)=i\}}m_{td}[v]\cdot m_{bu}[v.j]$ (3)
Projected Gradient

### v2

# FourierCSP: Differentiable Constraint Satisfaction Problem Solving by Walsh-Fourier Expansion

@@ Abstract
The Constraint-Satisfaction Problem (CSP) is fundamental in mathematics, physics, and theoretical computer science. Continuous Local Search (CLS) solvers, as recent advancements, can achieve highly competitive results on certain classes of Boolean Satisfiability (SAT) problems. Motivated by these advances, we extend the CLS framework from Boolean SAT to general CSP with finite-domain variables and expressive constraint formulations. We present FourierCSP, a continuous optimization framework that generalizes the Walsh-Fourier transform to CSP, allowing for transforming versatile constraints to compact multilinear polynomials, thereby avoiding the need for auxiliary variables and memory-intensive encodings. We employ projected subgradient and mirror descent algorithms with provable convergence guarantees, and further combine them to accelerate gradient-based optimization. Empirical results on benchmark suites demonstrate that FourierCSP is scalable and competitive, significantly broadening the class of problems that can be efficiently solved by differentiable CLS techniques and paving the way toward end-to-end neurosymbolic integration.
Machine Learning, ICML
1 Introduction
The Constraint-Satisfaction Problem (CSP) is foundational in mathematics, physics, and theoretical computer science. Formally, a CSP consists of variables defined over finite domains and a set of constraints restricting their values; the objective is to identify an assignment that satisfies all constraints Applications of CSP solving span diverse fields, including artificial intelligence (Baluta et al., 2019), database query (Kolaitis and Vardi, 1998), operational research (Brailsford et al., 1999), statistical physics (Nagy et al., 2024), information theory (Golia et al., 2022), and quantum computing (Vardi and Zhang, 2022). Despite their significant importance, CSP inherently exhibits a combinatorial nature and is generally NP-complete, presenting a substantial challenge in solving it efficiently. Consequently, extensive research and development efforts have been dedicated to advancing CSP solving (Schulte et al., 2006; Hebrard, 2008; Kuchcinski and Szymanek, 2013; Prud’homme and Fages, 2022; Perron et al., 2023; Krupke et al., 2024).
Although classical CSP engines (Schulte et al., 2006; Hebrard, 2008; Kuchcinski and Szymanek, 2013; Prud’homme and Fages, 2022) are powerful, their reliance on backtracking and discrete propagation creates a fundamental barrier for deep learning integration. The lack of differentiability precludes backpropagation, hindering the embedding of constraint reasoning into neural architectures. Recent hybrid approaches have attempted to bridge this gap by learning heuristics for branching and cut selection (Khalil et al., 2016; Zhang et al., 2022; Ling et al., 2024; Vo et al., 2025; Zeng et al., 2025; Gasse et al., 2019). However, these methods merely augment the discrete search process and do not yield differentiable solvers.
In parallel, Continuous Local Search (CLS) approaches for Boolean Satisfiability (SAT) have demonstrated that gradient-based optimization on smooth multilinear relaxations can yield highly effective differentiable solvers (Kyrillidis et al., 2020, 2021b). Yet, these formulations rely on the algebraic properties of Boolean variables $\{\pm 1\}$, making them inapplicable to the multi-valued domains and rich constraints found in general CSPs. Existing differentiable CSP techniques generally rely on ad-hoc heuristics or specific constraint types (Xu et al., 2018; Goyal et al., 2024; Xu et al., 2025), lacking a unified formulation for general finite-domain constraints.
This gap necessitates a differentiable framework that: (i) operates directly on finite-domain variables, avoiding the dimensionality explosion of Boolean encodings; (ii) can formulate versatile constraints; and (iii) ensures the continuous relaxation faithfully captures the structure of the discrete feasibility region. To achieve this, we utilize the Walsh-Fourier basis, which provides an orthogonal multilinear expansion that yields smooth optimization landscapes, ideal for projected subgradient and mirror-descent methods. While Fourier-based relaxations have proven effective for Boolean SAT (Kyrillidis et al., 2020, 2021b), in this work, we generalize the Walsh-Fourier transform to the finite-domain setting. This provides a principled algebraic foundation for differentiable CSP solving, extending the capabilities of CLS beyond the Boolean domain. The contribution of this paper can be summarized as follows:
A Uniform Differentiable CSP Framework: We provide a rigorous algebraic generalization of the Walsh-Fourier transform for variables with finite domains. The proposed framework is called FourierCSP. It established a principled basis for differentiable constraint solving.
Provable Optimization Algorithm: In addition to the FourierCSP framework, we develop specialized variants of projected subgradient and mirror-descent methods tailored to this setting. We provide rigorous convergence guarantees, ensuring the reliability of the continuous search.
Efficient Implementation and Empirical Evaluation: We conduct extensive evaluations on standard benchmark suites. Our results demonstrate that FourierCSP is highly scalable and competitive with existing baselines, significantly broadening the scope of problems solvable by differentiable techniques.
2 Preliminaries
@@
The Walsh-Fourier Transform is a method for transforming a function with a finite domain into a multilinear polynomial. Given the Fourier bases, any function defined on a finite domain has an equivalent Fourier representation.
Definition 2.1 (Fourier Bases).
@@
Theorem 2.2 (Walsh-Fourier Expansion (O’Donnell, 2014)).
@@
3 Theoretical Framework
In this section, we first recap the previous CLS for Boolean SAT solving. We then demonstrate how our proposed solver, FourierCSP, can extend the CLS framework to general constraint satisfaction problems. Proofs are delayed to Appendix A.
3.1 Continuous Local Search
[…]
Definition 3.1 (Objective).
Given a formula $F=\bigwedge_{c\in{}C}c$, every constraint $c\in{}C$ can be transformed to a function by Theorem 2.2. We denote the corresponding expansion as $f_{c}$ and the objective function w.r.t. $F$ by
$\mathcal{F}=-\sum_{c\in{}C}f_{c}.$
@@ 3.1 Continuous Local Search
The entire lineage of CLS solvers, from the 1990s approaches (Gu, 1994) to modern variants, is exclusively designed for Boolean constraints. While they perform well on certain problem classes, these solvers often fail to scale effectively to large or complex CSP instances. This limitation becomes particularly acute for finite-domain problems, which require conversion to potentially larger Boolean formulas. This limitation fundamentally restricts their practical application for real-world problems that naturally require finite-domain representations.
3.2 FourierCSP
To address the above limitation, we propose FourierCSP, a CLS framework that directly supports finite-domain CSPs without requiring Booleanization. We start by applying randomized rounding to the generalized form of WFE in Theorem 2.2, which reduces the CSP to a differentiable multilinear function over a product of probability simplices. In particular, although stationary points may occur in the interior of the domain, we prove that every local minimum of the multilinear objective must lie on the boundary of the simplices, ensuring the solution quality. Building on this differentiable formulation, we analyze the behavior of Projected Gradient Descent (PGD) and Mirror Descent (MD) and derive practical convergence bounds for both methods. Based on these theoretical insights, we propose a Hybrid Descent (HD) that combines PGD and MD, exploiting their complementary geometries to accelerate convergence.
@@ 3.2 FourierCSP
Theorem 2.2 illustrates how to expand discrete functions to multilinear polynomials through WFE. The resulting Fourier coefficients $\hat{f}(\alpha)$ provide valuable insights into analyzing the discrete function properties. However, the expanded function $f$ remains discrete and non-differentiable given the Fourier Bases in Definition 2.1, limiting its utility for gradient-based optimization. To enable continuous optimization, we introduce a randomized rounding $\mathcal{R}$ that relaxes discrete variable assignments into a continuous domain. Consider a set of finite domain variables $X=\{x_{1},...,x_{n}\}$. The vector probability space of $x_{i}$ is a simplex denoted by
@@
Definition 3.2 (Randomized Rounding).
@@
Definition 3.3 (Probability Space).
We define a probability space on a discrete matrix w.r.t. to real point $P\in\tilde{D}$, denoted by $\mathcal{S}:D\to[0,1]$, as:
@@
The probability space allows us to define a continuous extension of any discrete function by taking its expectation over a randomized assignment.
Theorem 3.4 (Expectation).
@@
$\mathop{\mathbb{E}}_{X\sim\mathcal{S}_{P}}[\mathcal{F}(X)]$ $=\mathop{\mathbb{E}}_{X\sim\mathcal{S}_{P}}[-\sum_{c\in{}C}f_{c}(X)]$ (1)
$=-\sum_{c\in{}C}\sum_{\alpha}\left(\hat{f}_{c}(\alpha)\cdot\prod_{i=1}^{n}p_{i,\alpha_{i}}\right)$
$=\mathcal{F}(P).$
We can observe from Theorem 3.4 that $\mathcal{F}(P)$ is a multilinear polynomial. Therefore, we relax the domain from discrete $D$ to continuous $\tilde{D}$, and it leads to a $\rho$-Lipschitz, and $L$-smooth multilinear continuous function.
Proposition 3.5 (Lipschitz).
Let $\mathcal{F}:\tilde{D}\to[-|C|,0]$ be a multilinear polynomial. Let $n$ be the number of variable, then $N=\sum_{i=1}^{n}|\mathrm{dom}(x_{i})|$. For every point $P,P^{\prime}\in\tilde{D}$, we have
$|\mathcal{F}(P)-\mathcal{F}(P^{\prime})|$ $\leq\rho\|P-P^{\prime}\|,$
$|\nabla\mathcal{F}(P)-\nabla\mathcal{F}(P^{\prime})|$ $\leq L\|P-P^{\prime}\|,$
where $\rho=|C|N^{\frac{1}{2}}$, $L=|C|N$.
As the expectation is a continuous and smooth multilinear polynomial, we formalize the reduction in Theorem 3.6.
Theorem 3.6 (Reduction).
@@
$\min_{P\in\tilde{D}}\left(\mathop{\mathbb{E}}_{X\sim\mathcal{S}_{P}}[\mathcal{F}(X)]\right)=-|C|.$ (2)
3.3 Optimization Approaches
The essence of CLS is using continuous optimization to find the minima of a non-convex but smooth function, where the smoothness of the WFE was analyzed in Proposition 3.5. Given the inherent difficulty of optimizing non-convex functions (Jain and Kar, 2017), it is often sufficient in practice to converge to a local optimum and subsequently verify, via Theorem 3.6, whether it corresponds to a global solution. During continuous optimization, the descent steps move the assignment toward regions of lower objective value, and the iterate eventually reaches a local minimum at which the suboptimality gap $\mathcal{F}(P_{t})-\mathcal{F}(P^{*})$ becomes small, typically $\leq\epsilon$. A potential concern is the existence of local minima strictly inside the feasible region, i.e., in $\tilde{D}\setminus\delta\tilde{D}$, since such points would provide no guarantee about the quality of the decoded discrete assignment. However, Lemma 3.7 establishes that no interior local maxima exist for our multilinear objective, and thus all optimal points lie on the boundary. This structural property directly leads to Theorem 3.8.
Lemma 3.7.
Let $\tilde{D}_{\perp}$ be a hyperplane with $\vec{1}$ as the normal vector. For every non-constant $\mathcal{F}$ and a point $P\in\tilde{D}\backslash{\delta\tilde{D}}$, there exist a region size $\epsilon>0$ and a directions, $V\in\tilde{D}_{\perp}$ such that for all $\eta\in(0,\epsilon)$, the following holds.
$\mathcal{F}(P-\eta{}V)<\mathcal{F}(P)$
Theorem 3.8 (Optimality).
Any interior point $P\in\tilde{D}\backslash{}\delta\tilde{D}$ cannot be a local optimum of $\mathcal{F}(P)$. The optimal points are only attained from the boundary $\delta\tilde{D}$.
Having established these optimization characteristics, the next question is how different descent geometries behave within this landscape. In particular, PGD and MD offer distinct update rules and induce different optimization geometries on the probability simplices.
Projected Gradient Descent